archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details We have an issue on the SFR devices in Département 83 A ticket has been opened A fibre has been cut Date Tuesday 03 December

    Original URL path: http://status.ovh.es/?do=details&id=5886 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    the 59 Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details A incident was detected on SFR equipment in the department 59 A ticket has been opened Date Sunday 01 December 2013 20

    Original URL path: http://status.ovh.es/?do=details&id=5883 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    2013 09 16AM Reason for closing Done Comments 4 Comment by OVH Sunday 01 December 2013 20 37PM The site BEA13 is not accessible anymore Marseille loop is no longer secure Sites other than BEA13 have no cut off Comment by OVH Sunday 01 December 2013 20 37PM We detected a fault in the optical equipment we will intervene on site Comment by OVH Monday 02 December 2013 09 15AM

    Original URL path: http://status.ovh.es/?do=details&id=5881 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    5820 FS 9735 rsa59 Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details The DSLAM crashed it s now up again Date Wednesday 20 November 2013 16 58PM Reason for closing Done Comments

    Original URL path: http://status.ovh.es/?do=details&id=5820 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    to resolve the management problem on the DSLAM No impact on clients synchronisation Date Wednesday 08 October 2014 15 14PM Reason for closing Done Comments 5 Comment by OVH Tuesday 19 November 2013 19 00PM The switchover has not resolved the problem we will reboot the DSLAM tonight at around 1 00 outage of 10min expected Comment by OVH Wednesday 20 November 2013 02 28AM saving conf Comment by OVH

    Original URL path: http://status.ovh.es/?do=details&id=5805 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details Following works on the fiber of Lille loop the DSLAM of LQN59 is currently inaccessible It will be back running no later than 6 00 Date Friday 15 November 2013 06 50AM Reason for

    Original URL path: http://status.ovh.es/?do=details&id=5782 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details An incident has been detected on the France Telecom equipment involving a sharp cut in the DID26 It should be fixed by 2 30pm today Date Wednesday 30 October 2013 16 19PM Reason for closing Done Comments 1

    Original URL path: http://status.ovh.es/?do=details&id=5696 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Type Incident Category Backend Core Status CLOSED Percent Complete Details An incident has been detected on the SFR equipment in the departement 31 A ticket is open Date Thursday 31 October 2013 11 48AM Reason for closing Done Comments 1

    Original URL path: http://status.ovh.es/?do=details&id=5695 (2015-10-15)
    Open archived version from archive



  •